OpenStudy (anonymous):

Write an equation of a line with a slope of 0 and passing through the point (5,4). Hint: To answer this question, consider the type of line that has a slope of 0. Then write an equation of that line using the values in the ordered pair.

OpenStudy (anonymous):

horizontal line = zero slope

OpenStudy (cwrw238):

line with a slope of zero is horizontal so it passes through y = 4 and is parallel to x-axis 

OpenStudy (cwrw238):

that should be 4 in diagram - not 5

OpenStudy (anonymous):

so what will the equation be

OpenStudy (anonymous):

y = a(x) + b where a = 0

OpenStudy (anonymous):

what bout x and b?

OpenStudy (anonymous):

y=0(x)+b

OpenStudy (anonymous):

x is wiped out because of zero

OpenStudy (anonymous):

you just need to find b

OpenStudy (anonymous):

can u help me please

OpenStudy (anonymous):

All x values have the same y value

OpenStudy (anonymous):

b is the y intersect

OpenStudy (anonymous):

i need help im so confused now

OpenStudy (anonymous):

so b is 4

OpenStudy (anonymous):

y=0(x)+4 ?

OpenStudy (anonymous):

hello?